Dear Tutor,
Thank you for the time you are taking to answer my question.
The question is: Kagiso jogs arund a rectangular track with a perimeter of 850m. He measured one side that worked out to be 105m. he jogs around this track 7 times. a)how far does he run every morning and B) how long will the other side of the track be then.
my own workings was that you would multiply 105x7 to find out how far he ran and that the other side would also be 105.
There would still be two sides left of 640 and if you divided it into two it would work out to two sides of 105 and two sides of 320.
Please let me know if I have worked this out correctly and if not, please help if possible.
thank you again
Tracey

Answer by Maths68(1474)

Perimeter of rectangle joging track = 850m
One side = 105m
He jogs around the track= 7 times
We need to know:-
a)how far does he run every morning
He jogs = 850 * 7 =5950m = 5950/1000 = 5.95km
B) how long will the other side of the track be

Perimeter of rectangle = 2(one side + other side)
850 = 2(105+other side)
850 = 210+2(other side)
850-210 = 2(other side)
640 = 2(other side)
640/2=2(other side)/2
320=other sides
other side=320m